OpenStudy (anonymous):

Find the indicated output for this function. if h(x)=x^3-2, find h(2) @mathstudent55 @heather040200 @Sceenie

OpenStudy (anonymous):

h sub x .. sub2 for x & solve

OpenStudy (anonymous):

it will look like h(2)=2^3-2 2 * 2 * 2 is 8 h(2)=8-2 h(2)=6 2h=6 2/2 = 6/2 h=3

OpenStudy (anonymous):

thanks @Sceenie

OpenStudy (anonymous):

youre welcome

OpenStudy (anonymous):

can i message you @Sceenie